if a cube has a surface area of 726 what would each side equal. this is very important. my grade for math depends on this question please help me asap. thankyou
Answer by jim_thompson5910(35256)

SA = 6s^2

726 = 6s^2

726/6 = s^2

121 = s^2

s^2 = 121

s = sqrt(121)

s = 11


So the length of each side would equal 11